Fermi National Laboratory


Accelerator Update

Monday March 15
The day shift began with the Tevatron (TeV) ten-day shutdown just beginning. Linac and Booster remain interlocked for special testing.

A Linac expert turned off the H- Source at 8:15 AM.

The Antiproton Source (Pbar) recovered from the planned feeder 24 power outage at 8:28 AM.

A RF expert took down the Main Injector (MI) and TeV RF at 8:37 AM.

At 9:15 AM, FESS technicians began cleaning the TeV pond strainers.

At 3:45 PM, Operations received a fire alarm for the Preaccelerator. Operators responded and met the Fire Department. They saw smoke in the H- Source area, but no fire. Two Linac experts investigated further and believed the smoke was from a slipping belt for a vacuum pump.

A power supply tech repaired the Meson Center power supply (MC1D) that knocked the MCenter experiment off over the weekend. It was a blown LCW hose. The power supply must dry out before conducting any operational tests.

At 5:32 PM, operators manually turned off the MI60 sector pond water pumps; remote control wasn't working. Pond water was leaking around the strainers into the service building because two pumps were running at the same time. Typically, only one pump runs at a time. Operators called in an expert to help straighten out the problem.

Tuesday March 16
The midnight shift began with the second day of shutdown.

Cryo system techs began work on the TeV sector D1 wet engine at 6:42 AM.

Operations made Linac ready for access at 8 AM.

Operators finished interlocking all the Pbar tunnels and transfer lines at 5:30 PM for safety system tests.

Operators finished interlocking all the TeV enclosures at 8:49 PM for safety system tests.

Wednesday March 17
The midnight shift began with the third day of shutdown.

At 6:40 AM, cryo system techs reported beginning work on various wet engines around the TeV ring.

A Partial List of Work Completed or Started

Linac

  • Installed 20 MHz digitizers
  • Pulled loss monitor cables
  • Installed loss monitors
  • Replaced filament controller for KRF1

Controls

  • Want to restart the Java nodes today to install new code; it should take about 2.5 hours
  • Troubleshoot a power supply UPS problem in MI

Booster

  • Installed vacuum control racks
  • Replaced the Dog13 power supply
  • Worked on kickers
  • Fixed a flow switch

Main Injector

  • Completed the quadrupole magnet roles in the P1 line

Recycler

  • Installed loss monitor
  • Completed cable pulls in 62
  • Completed the re-buss for the Lambertson 52 magnet
  • Think they have found alignment problems for e-cool

TeV

  • The cryo system at A4 should be warm later today
  • The P-150 moves were completed
  • Completed the magnet roll measurements at sectors E and 1/2 of B
  • Continue work on the truck/quadrupole resonant noise problem

Cryo

  • The A4 cryo system warm-up will finish later today
  • Reported that they will have completed 17 wet engine overhauls by the end of the shutdown

Instrumentation

  • They have fixed the TeV sector E1 flying wire work as well as the Pbar flying wire work
  • Working on the MI DC current monitor
  • Working on the MI8 BPM phase match problem

Pbar

  • They have completed seven of their forty-one jobs

RF

  • Conducting TeV and MI maintenance

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 3/17/2004   email Fermilab